THE PROBABILITIES THAT A SURGEON OPERATES ON 3,4,5,6 or patients in any one day are 0.15,0.20,0.25,0.20 and 0.20 respectively


1
Expert's answer
2022-02-22T04:06:48-0500

The number of patients are {3,4,5,6,7} occurring with the probabilities {0.15,0.20,0.25,0.2,0.2}


The Mean number of Patients is

"=3(0.15)+4(0.20)+5(0.25)+6(0.2)+7(0.2)\\=0.45+0.8+1.25+1.2+1.4\\=5.1"